### v1

# On Nontrivial Winning and Losing Parameters of Schmidt Games

Abstract.
In this paper we completely describe the winning and losing conditions different from the only “trivial” conditions known before. In other words, we solve the open question of finding a complete nontrivial Schmidt diagram. In addition, we give the new bounds for two family of sets: one related to frequencies of digits in base-$2$ expansions, and one connected to the set of the badly approximable numbers.
1. Set-up and Introduction: Schmidt Games
Wolfgang Schmidt first introduced Schmidt games in [S66] in order to study the set of badly approximable numbers. Winning sets in these games have special properties, such as a full Hausdorff dimension; as such, Schmidt games have been used to prove results in number theory and dynamical systems. Here, we briefly describe the setup of a Schmidt $(\alpha,\beta)$-game. Throughout this paper we work with the Schmidt game on $\mathbb{R}$, so for simplicity we give the definitions only in this special case.
Denote by $I$ the open unit square:
$I\overset{\rm{def}}{=}\{(\alpha,\beta):0<\alpha,\beta<1\}=(0,1)\times(0,1)\,.$
Then, fix a set $S\subset{\mathbb{R}}$ and $(\alpha,\beta)\in I$. Two players, Alice and Bob, can play a Schmidt game on $S$, taking alternating turns, as follows: Bob begins by choosing an interval $B_{0}$. Given an interval $B_{i}$, Alice chooses an interval $A_{i}\subset B_{i}$ such that $|A_{i}|=\alpha|B_{i}|$. Similarly, given an interval $A_{i}$, Bob chooses an interval $B_{i+1}\subset A_{i}$ such that $|B_{i+1}|=\beta|A_{i}|$. Let
$x_{\infty}\overset{\rm{def}}{=}\bigcap_{k=1}^{\infty}B_{k}=\bigcap_{k=1}^{\infty}A_{k}.$
We call $S$ $(\alpha,\beta)$-winning if, regardless of how Bob chooses his intervals, Alice can play in such a way that $x_{\infty}\in S$. For every $S$, define the Schmidt diagram $D(S)$ of $S$ as the set of all pairs $(\alpha,\beta)\in I$ such that $S$ is $(\alpha,\beta)$-winning. In his original paper, Schmidt proved that there are two “trivial zones” of Schmidt diagrams:
Lemma 1.1.
[S66]
  • (a)
    If $S$ is dense, then $D(S)$ contains the set
    $\check{D}:=\big{\{}(\alpha,\beta)\in I:\beta\geq(2-\alpha)^{-1}\big{\}},$
    with equality if $S$ is countable (see Figure 2).
  • (b)
    If $S\subsetneq{\mathbb{R}}$, then $D(S)$ is contained in the set
    $\hat{D}:=\big{\{}(\alpha,\beta)\in I:\beta>2-\alpha^{-1}\big{\}},$
    with equality if the complement of $S$ is countable (see Figure 2).
Figure 1. $\check{D}$ (blue)
Figure 1. $\check{D}$ (blue)
Figure 2. $\hat{D}$ (blue)
There are only four sets which were known to coincide with $D(S)$ for some $S$: $\emptyset$, $\check{D}$, $\hat{D}$, and $I$. Our goal with this project was to find some set $S$ such that $\check{D}\subsetneq D(S)\subsetneq\hat{D}$ and to completely describe $D(S)$. It is worth noting that this problem was investigated in an earlier PRIMES paper, which investigated sets relating to particular decimal expansions [KZT12]. In the aforementioned paper, some nontrivial statements about Schmidt diagrams were obtained, but no $D(S)$ different from the four diagrams discussed above was completely described.
2. Main Results
One of the main results of this paper is a complete description of a Schmidt Diagram different from $\emptyset$, $\check{D}$, $\hat{D}$ and $I$. More specifically, we prove the following theorem.
Theorem 2.1.
There exists a set $\mathscr{S}$ such that $D(\mathscr{S})=\{(\alpha,\beta):0<\alpha<\beta<1\}$ (see Figure 3).
Figure 3. $D(\mathscr{S})$ (blue)
We develop the techniques for this in Section 3, and the more general statements about sets with such a diagram are given in Theorem 3.10 and Corollary 3.11. We construct a special family of examples in Section 4.
In Section 5, we describe some winning and losing conditions for families of sets connected with digit frequencies in the dyadic expansion of real numbers. We prove a general case in Theorem 5.3 and a special symmetric case in Corollary 5.6. In Section 6 we give some new results about winning and losing conditions for families of sets connected with badly approximable numbers in base $B$ expansion: specifically in Theorems 6.4, 6.5, and 6.6.
The bounds for the winning and losing zones we describe in Sections 5 and 6 are not optimal; however, they give some estimates and shed some light on how their respective Schmidt diagrams can look like. Figures 5 and 5 show two examples of the possible bounds (where the blue zones are subsets of the corresponding Schmidt diagram and the red zones are subsets of its complement). The sets $D_{c}^{-}$ used in the first example will be defined later in Section 5 and the sets $\operatorname{\widetilde{2-BA}}(c)$ will be defined in Section 6.
Figure 4. Bounds for $D(D_{c}^{-})$ with $c=0.7$
Figure 4. Bounds for $D(D_{c}^{-})$ with $c=0.7$
Figure 5. Bounds for $D(\operatorname{\widetilde{2-BA}}(c))$ with $c=0.035$
3. General strategies in the Schmidt’s game
This section gives some general statements concerning strategies in Schmidt’s game.
Let us start with a classical result due to Schmidt:
Lemma 3.1.
[S66] If a set $S$ is $(\alpha,\beta)$-winning, $\alpha\beta=\alpha^{\prime}\beta^{\prime}$, and $\alpha>\alpha^{\prime}$, then $S$ is also $(\alpha^{\prime},\beta^{\prime})$-winning.
Sometimes for simplicity we will also refer to existence the of positional strategies, i.e. to the following classical statement:
Theorem 3.2.
[S66] If a set $S$ is winning, there exists a positional winning strategy; that is, a strategy that only depends on the most recent interval played and not the prior history of moves.
The next four statements are the technical steps in proving Lemma 3.7; however, they are rather general and can be used separately in the further research.
Lemma 3.3.
Fix $m\in\mathbb{N}$. Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$. Suppose Bob has a strategy for $(\alpha,\beta)$-game with the set $D$ that allows him to win, and the first ball in this strategy is $B_{0}$. Then:
  1. (1)
    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $D$ starting with the ball $m^{k}B_{0}$;
  2. (2)
    For any $k\in\mathbb{Z}$ Bob has a winning strategy for $(\alpha,\beta)$-game with the set $D$ starting with the ball $m^{k}+B_{0}$.
(Here $m^{k}B_{0}=\{m^{k}x:\,\,\,x\in B_{0}\}$ and $m^{k}+B_{0}=\{m^{k}+x:\,\,\,x\in B_{0}\}$).
Proof.
  1. (1)
    Suppose Bob starts the game with $B_{0}^{\prime}=m^{k}B_{0}$, and Alice plays some ball $A_{0}^{\prime}\subset B_{0}^{\prime}$. Let us define $A_{0}=m^{-k}A_{0}^{\prime}\subseteq B_{0}$. Now let $B_{1}\subset A_{0}$ be the ball that Bob would play after $A_{0}$ according to his winning strategy. Let Bob play $B_{1}^{\prime}=m^{k}B_{1}\subset A_{0}^{\prime}$.
    After $n$ steps, analogously, suppose Alice plays some ball $A_{n}^{\prime}\subset B_{n}^{\prime}$. Define $A_{n}=m^{-k}A_{n}^{\prime}$, and define $B_{n+1}$ as the ball Bob would play after $A_{n}$; then let Bob play $B_{n+1}^{\prime}=m^{k}B_{n+1}\subset A_{n}^{\prime}$.
    It is easy to see that $\bigcap\limits_{i=0}^{\infty}B_{i}^{\prime}=m^{k}\bigcap\limits_{i=0}^{\infty}B_{i}=m^{k}x$, where $x$ is the outcome of some game with the starting ball $B_{0}$, i.e. $x\notin D$, therefore $m^{k}x\notin D$ and Bob wins.
  2. (2)
    The strategy and the argument is analogous; we just take $A_{i}=A_{i}^{\prime}-m^{k}$ and $B_{i}^{\prime}=B_{i}+m^{k}$ and use the fact that $x\in D$ if and only if $m^{k}+x\in D$.
Lemma 3.4.
Let $S\subseteq\mathbb{R}$ be any set. Suppose Bob has a strategy for $(\alpha,\beta)$-game with the set $S$ that allows him to win, and the first ball in this strategy is $B_{0}$. Then for any $(\alpha^{\prime},\beta^{\prime})$ such that $\beta^{\prime}<\beta$ and $\alpha^{\prime}\beta^{\prime}=\alpha\beta$ Bob has a winning strategy starting from $B_{0}$ in $(\alpha^{\prime},\beta^{\prime})$-game.
Proof.
Suppose Bob starts $(\alpha^{\prime},\beta^{\prime})$-game with the ball $B_{0}^{\prime}=B_{0}$, and Alice plays some ball $A_{0}^{\prime}$. Let $A_{0}$ be the ball of size $|A_{0}|=\alpha|B_{0}|=\frac{\alpha}{\alpha^{\prime}}|A_{0}^{\prime}|<|A_{0}^{\prime}|$ centered at the center of $A_{0}^{\prime}$; suppose $B_{1}$ is the ball Bob would play after $A_{0}$ in $(\alpha,\beta)$-game. Let Bob play the ball $B_{1}^{\prime}=B_{1}$ after $A_{0}^{\prime}$ in $(\alpha^{\prime},\beta^{\prime})$-game; indeed, $B_{1}^{\prime}\subset A_{0}\subseteq A_{0}^{\prime}$ and $|B_{1}^{\prime}|=\beta|A_{0}|=\alpha\beta|B_{0}^{\prime}|$, so Bob can make such a move.
After $n$ steps, analogously, suppose Alice plays some ball $A_{n}^{\prime}\subset B_{n}^{\prime}$. Define $A_{n}$ to be the ball of size $|A_{n}|=\alpha|B_{n}|=\frac{\alpha}{\alpha^{\prime}}|A_{n}^{\prime}|<|A_{n}^{\prime}|$ centered at the center of $A_{n}^{\prime}$, and define $B_{n+1}$ as the ball Bob would play after $A_{n}$; then let Bob play $B_{n+1}^{\prime}=B_{n+1}\subset A_{n}^{\prime}$. One has $\bigcap\limits_{i=0}^{\infty}B_{i}^{\prime}=\bigcap\limits_{i=0}^{\infty}B_{i}=x$ where $x$ is the outcome of some game with the starting ball $B_{0}$, i.e. $x\notin S$, therefore Bob wins.
Lemma 3.5.
Fix $m\in\mathbb{N}$. Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$. Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $D$. Then for any $L>0$ and any $\varepsilon>0$ there exists a winning strategy in this game for Bob starting with $B_{0}$ with
$L\leq|B_{0}|<(1+\varepsilon)L.$
Proof.
Suppose the ball Bob starts his winning strategy with has size $l$. We denote $\xi=-\log_{\alpha\beta}m$ and $\chi=\log_{\alpha\beta}(\frac{L}{l})$, and let $\varepsilon^{\prime}=\log_{\alpha\beta}(1+\varepsilon)$.
As $\xi$ is irrational, the sequence $k\xi+\chi+\varepsilon^{\prime}$ is uniformly distributed modulo 1, so there exists $k$ such that $0<\{k\xi+\chi+\varepsilon^{\prime}\}\leq\varepsilon^{\prime}$. Let $n=\lfloor k\xi+\chi+\varepsilon^{\prime}\rfloor$; in this case
$\log_{\alpha\beta}L+k\xi\leq n+\log_{\alpha\beta}l<\log_{\alpha\beta}L+k\xi+\varepsilon^{\prime}$
and
$\frac{L}{m^{k}}\leq(\alpha\beta)^{n}l=|B_{n}|<(1+\varepsilon)\frac{L}{m^{k}},$
where $B_{n}$ is any ball Bob can play using his strategy. Fix some possible $B_{n}$; if it occurs in the game, Bob can start the game with this $B_{n}$. By Lemma 3.3 Bob can start also with $B_{0}=m^{k}B_{n}$ and win. It remains to notice that $|B_{0}|=m^{k}|B_{n}|$ and
$L\leq|B_{0}|<(1+\varepsilon)L.$
Lemma 3.6.
Let $S\subseteq\mathbb{R}$ be any set. Suppose Bob has a strategy for $(\alpha,\beta)$-game with the set $S$ that allows him to win, and the first ball in this strategy is $B_{0}$. Then for any $\varepsilon>0$ and for any $B_{0}^{\prime}\subseteq B_{0}$ with $|B_{0}^{\prime}|\geq\frac{|B_{0}|}{1+\varepsilon}$, Bob has a winning strategy starting from $B_{0}^{\prime}$ in $((1+\varepsilon)\alpha,\frac{\beta}{1+\varepsilon})$-game.
Proof.
Fix $\varepsilon>0$. Suppose Bob starts $(\alpha^{\prime},\beta^{\prime})$-game (where $\alpha^{\prime}=(1+\varepsilon)\alpha$ and $\beta^{\prime}=\frac{\beta}{1+\varepsilon}$) with some $B_{0}^{\prime}$ such that $|B_{0}^{\prime}|=\frac{|B_{0}|}{1+\varepsilon}$.
Suppose Alice chooses some $A_{0}^{\prime}\subseteq B_{0}^{\prime}\subseteq B_{0}$; one has $|A_{0}^{\prime}|=\alpha^{\prime}|B_{0}^{\prime}|=(1+\varepsilon)\alpha\frac{|B_{0}|}{1+\varepsilon}=\alpha|B_{0}|$. Let $A_{0}=A_{0}^{\prime}$, and let $B_{1}$ be the ball Bob would play after $A_{0}$ in $(\alpha,\beta)$-game. Now, define $B_{1}^{\prime}$ as a ball centered at the center of $B_{1}$ having $|B_{1}^{\prime}|=\frac{|B_{1}|}{1+\varepsilon}$, and let Bob play $B_{1}^{\prime}$ after $A_{0}^{\prime}$ in our $(\alpha^{\prime},\beta^{\prime})$-game.
Analogously, let $A_{n}=A_{n}^{\prime}$, $B_{n+1}$ the ball Bob would play after $A_{n}$ in $(\alpha,\beta)$-game and $B_{n+1}^{\prime}$ a ball centered at the center of $B_{n+1}$ having $|B_{n+1}^{\prime}|=\frac{|B_{n+1}|}{1+\varepsilon}$; Bob can play $B_{n+1}^{\prime}$ after any possible $A_{n}^{\prime}$.
One can see that $\bigcap\limits_{i=0}^{\infty}B_{i}^{\prime}=\bigcap\limits_{i=0}^{\infty}B_{i}=x$ where $x\notin S$, as it is the outcome of an $(\alpha,\beta)$-game where Bob followed his winning strategy. Therefore Bob wins.
Finally, for an arbitrary $|B_{0}^{\prime}|\geq\frac{|B_{0}|}{1+\varepsilon}$ let us note that $|B_{0}^{\prime}|=\frac{|B_{0}|}{1+\varepsilon^{\prime}}$ for $\varepsilon^{\prime}\leq\varepsilon$, which means that Bob can win $((1+\varepsilon^{\prime})\alpha,\frac{\beta}{1+\varepsilon^{\prime}})$-game on $S$ starting from $B_{0}^{\prime}$. Lemma 3.4 completes the proof. ∎
Lemma 3.7.
Fix $m\in\mathbb{N}$. Let $D\subseteq\mathbb{R}$ be some set with a property that $m^{k}D=m^{k}+D=D$ for any integer $k$. Let parameters $(\alpha,\beta)$ be such that $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, and suppose Bob wins $(\alpha,\beta)$-game with $D$. Then for any $\beta^{\prime}<\beta,\,\alpha^{\prime}=\frac{\alpha\beta}{\beta^{\prime}}$, and any closed ball $T$ Bob can win $(\alpha^{\prime},\beta^{\prime})$-game starting with any ball $T$ as his initial ball.
Proof.
Fix an arbitrary ball $T$. Also, fix $\varepsilon>0$, and let $L=|T|$. By Lemma 3.5 Bob has a winning strategy in $(\alpha,\beta)$-game with a starting ball $B_{0}$ s.t. $(1+\frac{\varepsilon}{2})L\leq|B_{0}|<(1+\varepsilon)L$. Take $\varepsilon^{\prime}<\varepsilon$ such that $|B_{0}|=(1+\varepsilon^{\prime})L$.
Now, let $k$ be such that $\frac{1}{m^{k}}<\frac{1}{2}\varepsilon^{\prime}L=\frac{1}{2}(|B_{0}|-|T|)$; let $b_{0}$ be the center of $B_{0}$ and $t$ the center of $T$. There exists $q\in\mathbb{Z}$ such that $|\frac{q}{m^{k}}+b_{0}-t|<\frac{1}{m^{k}}$. Consider the ball $B_{0}^{\prime}=B_{0}+\frac{q}{m^{k}}$: by Lemma 3.3, Bob has a winning strategy in $(\alpha,\beta)$-game with a starting ball $B_{0}^{\prime}$. In addition, note that $T\subseteq B_{0}^{\prime}$ and $|T|=\frac{|B_{0}^{\prime}|}{1+\varepsilon^{\prime}}$. By Lemma 3.6, Bob has a winning strategy in $((1+\varepsilon)\alpha,\frac{\beta}{1+\varepsilon})$-game with a starting ball $T$.
Now, for $(\alpha^{\prime},\beta^{\prime})$ as in Lemma it remains to take $\varepsilon$ under the condition $(1+\varepsilon)\alpha=\alpha^{\prime}$. ∎
The discussion above implies the following statement:
Lemma 3.8.
Let $m\in\mathbb{N}$ and $D,Q\subseteq\mathbb{R}$ be some sets with the following properties:
  • $m^{k}D=m^{k}+D=D$ for any integer $k$.
  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$.
Then:
  1. (1)
    If $\alpha\geq\beta$, then the set $D$ is $(\alpha,\beta)$-losing.
  2. (2)
    If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $D\sqcup Q$ is $(\alpha,\beta)$-winning.
Proof.
  1. (1)
    Suppose $\alpha\geq\beta$ and $D$ is $(\alpha,\beta)$-winning. In this case Alice has a positional winning strategy described by the function $A(I)$ (where $I\subset\mathbb{R}$ is a closed interval) for this game. Consider $(\beta,\alpha)$-game on the set $D$ and let Bob play the following strategy:
    • Bob chooses $B_{0}=A([0,1])$.
    • If Alice plays the interval $A_{n}$ at the $n$-th step, Bob plays $B_{n+1}=-A(-A_{n})$.
    Suppose $x=\bigcap\limits_{i=0}^{\infty}B_{i}$ is the outcome of the game. Then $-x=\bigcap\limits_{i=0}^{\infty}A(-A_{n})$ is the outcome of the $(\alpha,\beta)$-game where Bob plays the balls $[-1;0];-A_{0},-A_{1},\ldots$ and Alice plays using her winning strategy $A(I)$. Therefore, $-x\in D$, which means $-x\notin D$ and Bob wins.
    It means that the point $(\alpha,\beta)$ is winning and the point $(\beta,\alpha)$ is losing for the set $D$, which contradicts Lemma 3.1. Thus, the point $(\alpha,\beta)$ is losing.
  2. (2)
    The idea of the proof is the same as in (a). Suppose $D\sqcup Q$ is $(\alpha,\beta)$-losing; Take $\varepsilon>0$ under the condition $(1+\varepsilon)^{2}<\frac{\beta}{\alpha}$. Then by Lemma 3.7 for any closed ball $T$ Bob has a winning strategy for $\left(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)}\right)$-game starting from $T$. We use this statement to construct a $\left(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon)\right)$-winning strategy for Alice.
    Suppose Bob starts some $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$-game with the ball $B_{0}$. Then Alice picks some ball $A_{0}\subseteq B_{0}$ without any restrictions. If Bob plays the ball $B_{n}$ on $n$-th step, Alice responds by playing $A_{n}$ where $-A_{n}$ is the ball Bob would play in $(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)})$-game which started from $-A_{0}$ after the balls $-A_{0},-B_{1},\ldots,-B_{n}$ (in this game $-A_{i}$ are Bob’s balls and $-B_{i}$ are Alice’s). Let $x$ be the outcome of this $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$-game; then $-x$ is the outcome of $(\alpha(1+\varepsilon),\frac{\beta}{(1+\varepsilon)})$-game Bob played with his winning strategy, i.e. $-x\in(-D)$ and $x\in D\subseteq D\sqcup Q$. It shows that $(\frac{\beta}{(1+\varepsilon)},\alpha(1+\varepsilon))$ is a winning point for $D\sqcup Q$.
    But one has $\alpha<\alpha(1+\varepsilon))<\frac{\beta}{(1+\varepsilon)}$, which by Lemma 3.1 contradicts our initial assumption. Therefore $D\sqcup Q$ is $(\alpha,\beta)$-winning.
Now, the goal of the next part is to find the condition under which we can remove dyadic rationals from the set $D$ so that it remains winning for the parameters above.
We start with the following almost obvious statement:
Lemma 3.9.
Suppose $S$ is an $(\alpha,\beta)$-winning set and for some $\xi\in\mathbb{R}$ one has $S+\xi\subseteq D$. Then the set $D$ is $(\alpha,\beta)$-winning.
Proof.
If $A(I)$ was Alice’s winning positional strategy on the set $S$, then $A(I-\xi)+\xi$ is a winning positional strategy on the set $D$. ∎
Now, we use the lemmas we have to prove the following statement:
Theorem 3.10.
Fix $m\in\mathbb{Z}$. Let $D\subseteq\mathbb{R}$ be some set with the following properties:
  • $m^{k}D=m^{k}+D=D$ for any integer $k$.
  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$, where $Q\subseteq\mathbb{Q}$ is some set.
  • $D+\xi\subseteq D$ for some $\xi\notin\mathbb{Q}$.
Then:
  1. (1)
    If $\alpha\geq\beta$, then the set $D$ is $(\alpha,\beta)$-losing.
  2. (2)
    If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $D$ is $(\alpha,\beta)$-winning.
Proof.
  1. (1)
    follows from Lemma 3.8.
  2. (2)
    By Lemma 3.8 we know that the set $D\sqcup Q$ is $(\alpha,\beta)$-winning. As $Q\subseteq\mathbb{Q}$, one has $D\sqcup Q+\xi\subseteq D$ for any $\xi\notin\mathbb{Q}$. By Lemma 3.9 then, $D$ is $(\alpha,\beta)$-winning for $\beta>\alpha$.
Corollary 3.11.
Let $D\subseteq\mathbb{R}$ be some set with the following properties:
  • $m_{i}^{k}D=m_{i}^{k}+D=D$ for any integer $k$ and some $m_{1},\,m_{2}\in\mathbb{N}$ with property $\log_{m_{1}}(m_{2})\notin\mathbb{Q}$;
  • $D\sqcup(-D)\sqcup Q=\mathbb{R}$, where $Q\subseteq\mathbb{Q}$ is some set;
  • $D+\xi\subseteq D$ for some $\xi\notin\mathbb{Q}$.
If $\beta>\alpha$, then the set $D$ is $(\alpha,\beta)$-winning.
Indeed: for any parameters $(\alpha,\beta)$ in this case either $\log_{m_{1}}(\alpha\beta)$ or $\log_{m_{2}}(\alpha\beta)$ is irrational, so by Theorem 3.10 $D$ is $(\alpha,\beta)$-winning.
4. Example: sets with a nontrivial diagram
Theorem 3.10 and Corollary 3.11 give the complete description of the nontrivial Schmidt diagram for sets with special properties. The goal of this section is to give an example of a set satisfying these properties, thus proving Theorem 2.1.
Fix $\xi\notin\mathbb{Q}$. Then, we define an auxiliary set $\Xi$ by $\Xi=\{r+q\xi:\,\,r,q\in\mathbb{Q},\,q>0\}$. One can see that $-\Xi=\{r+q\xi:\,\,r,q\in\mathbb{Q},\,q<0\}$ is disjoint with $\Xi$, and one has $\Xi+\xi\subseteq\Xi$ by construction.
Now, let us focus on the set $E=\mathbb{R}\setminus\Big{(}\mathbb{Q}\cup\Xi\cup(-\Xi)\Big{)}$. We will divide it into the union of the sets $\overline{x}=\{r+q\xi+px:\,\,r,q,p\in\mathbb{Q},\,\,p\neq 0\}$ over $x\in E$. Easy to see that $\overline{x}$ is the equivalence class under relation $\sim$ defined as follows: $x\sim y$ if $y=r+q\xi+px$ with the restrictions as above, therefore $E$ is the disjoint union of the sets $\overline{x}$ over pairwise non-equivalent $x$.
Next, let $\underline{x}=\{r+q\xi+px:\,\,r,q,p\in\mathbb{Q},\,\,p>0\}$. Then $\overline{x}=\underline{x}\sqcup\underline{-x}$. Indeed: if $\underline{x}\,\text{and}\,\underline{-x}$ intersect, then $x\in\Xi\sqcup(-\Xi)$, which we do not consider. Finally, note that $\underline{-x}=-\underline{x}$.
Let us enumerate all the sets $\overline{x}$ by some index set $\mathscr{A}$ and denote the corresponding set $E_{\alpha}$ where $\alpha\in\mathscr{A}$. Choose some $x_{\alpha}\in A_{\alpha}$ and denote $\underline{x_{\alpha}}=C_{\alpha}$; then $C_{\alpha}\sqcup(-C_{\alpha})=E_{\alpha}$. Therefore for $C=\bigsqcup\limits_{\alpha\in\mathscr{A}}C_{\alpha}$ one has $C\sqcup(-C)=E$; by construction also $C+\xi_{j}=C$ (as this works for each of $\underline{x}$).
Finally, we define $\mathscr{S}=\mathscr{S}(\xi)=C\sqcup\Xi$. Then $-\mathscr{S}=(-C)\sqcup(-\Xi)$, and one has $\mathscr{S}\sqcup(-\mathscr{S})\sqcup\mathbb{Q}=\mathbb{R}$. One also has:
  • $m^{k}\mathscr{S}=m^{k}+\mathscr{S}=\mathscr{S}$ for any integer $k$ and any $m\in\mathbb{N}$;
  • $\mathscr{S}+\xi\subseteq\mathscr{S}$.
Therefore, by Theorem 3.10 and Corollary 3.11, $\mathscr{S}$ is $(\alpha,\beta)$-winning for $\beta>\alpha$ and $(\alpha,\beta)$-losing for $\beta\leq\alpha$. This proves Theorem 2.1.
5. Digit frequencies
Next, we will investigate a separate but similar construction regarding frequencies of digits in base-$2$ expansions of real numbers. It is known that these sets are closely related to density and dimension, and hence they have been studied in the past (see [V59, FFNS21]).
First, some setup. Consider the base-$2$ expansions of the form $x=x_{0}.x_{1}x_{2}\cdots$ where $x_{0}$ is an integer and $x_{i}\in\{0,1\}$ are the digits in the base-$2$ expansion of $x$. In this paper, if $x$ can be represented by two seperate expansions, we take the lexicographically minimal expansion (for example, $0.100\ldots$ would instead be written as $0.011\ldots$). We define
$d^{+}(x,j)=\limsup_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:x_{i}=j\}}{k}$
and similarly,
$d^{-}(x,j)=\liminf_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:x_{i}=j\}}{k}.$
Define the sets
$D^{+}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{+}(x,0)>c\},$
$D^{-}_{c}$ :=$:=\{x\in{\mathbb{R}}:d^{-}(x,0)>c\}.$
The following is known about the Schmidt diagrams of these sets.
Theorem 5.1.
[FFNS21] The set $D^{+}_{c}$ is winning for $\log(8\alpha)/\log(\alpha\beta)>c$ and losing for $\log(\alpha/8)/\log(\alpha\beta)\leq c$.
The results of our research expand on that bound. To do so, we require the following lemma, known as the Equidistribution Theorem:
Lemma 5.2.
[B09] If $\alpha$ is irrational, then the sequence $\alpha,2\alpha,3\alpha,\ldots$ is uniformly distributed modulo $1$.
Theorem 5.3.
Suppose $\log_{2}{\alpha\beta}\notin\mathbb{Q}$. Then, the sets $D^{\pm}_{c}$ are winning for $\alpha\leq\frac{1}{4}$ and $\log(4\alpha)/\log(\alpha\beta)>c$, and losing for $\beta\leq\frac{1}{4}$ and $\log(\alpha/4)/\log(\alpha\beta)\leq c$.
In the special case the bounds given by Theorem 5.3 are presented on Figure 5.
Proof.
We will show the winning strategy for Alice.
Suppose Bob begins by playing an interval with radius $r_{0}$. Alice will play in the following way: suppose Bob plays some interval $B_{t}$. Define $k_{t}=-\left\lfloor\log_{2}{(\alpha\beta)^{t}r_{0}}\right\rfloor$, and pick the smallest $m\in\mathbb{Z}$ such that $m/2^{k_{t}}\in B_{t}$. Then, it follows that the dyadic interval $\left(\frac{m}{2^{k_{t}}},\frac{m+1}{2^{k_{t}}}\right)\subset B_{t}$. Then, Alice plays $A_{t}=\left(\frac{m}{2^{k_{t}}},\frac{m}{2^{k_{t}}}+2\alpha(\alpha\beta)^{t}r_{0}\right)$. Define $l_{t}$ such that $k_{t}+l_{t}=-\left\lfloor\log_{2}{\alpha(\alpha\beta)^{t}r_{0}}\right\rfloor-2$. Then, Alice ensures that for any $x\in A_{t}$, the $(k_{t}+1)$th digit to the $(k_{t}+l_{t})$th digit is equal to 0. In other words, she guarantees $l_{t}$ zeroes on this turn. For convenience, let $a=\log_{2}\alpha$, $b=\log_{2}\beta$, and $\rho=\log_{2}r_{0}$; note that $a+b\notin\mathbb{Q}$. We have
$l_{t}$ $=(k_{t}+l_{t})-k_{t}$
$=\left\lfloor(a+b)t+\rho\right\rfloor-\left\lfloor a+(a+b)t+\rho\right\rfloor-2$
$=(a+b)t+\rho-(a+(a+b)t+\rho)-\{(a+b)t+\rho\}+\{a+(a+b)t+\rho\}-2$
$=-(a+2)+\{a+(a+b)t+\rho\}-\{(a+b)t+\rho\}.$
This implies that by the Equidistribution Theorem,
$\lim\limits_{n\rightarrow\infty}\frac{\sum_{i=0}^{n-1}l_{i}}{n}$ $=-(a+2).$
Now, notice that for $k_{n}\leq k<k_{n+1}$ one has
#$\frac{\#\{1\leq i\leq k:x_{i}=0\}}{k}=\frac{\#\{1\leq i\leq k_{n}:x_{i}=0\}}{k_{n}}\cdot\frac{k_{n}}{k}+\frac{\#\{k_{n}+1\leq i\leq k:x_{i}=0\}}{k},$
which implies
#$\frac{\#\{1\leq i\leq k_{n}:x_{i}=0\}}{k_{n}}\cdot\frac{k_{n}}{k_{n+1}}\leq\frac{\#\{1\leq i\leq k:x_{i}=0\}}{k}.$
It is easy to see that $k_{n+1}-k_{n}$ is bounded, therefore $\lim\limits_{n\rightarrow\infty}\frac{k_{n}}{k_{n+1}}=1$, and
$d^{-}(x,0)$ $=\liminf\limits_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:x_{i}=0\}}{k}$
$\geq\liminf\limits_{n\rightarrow\infty}\frac{\#\{1\leq i\leq k_{n}:x_{i}=0\}}{k_{n}}\cdot\frac{k_{n}}{k_{n+1}}$
$\geq\lim\limits_{n\rightarrow\infty}\frac{\sum_{i=0}^{n-1}(l_{i})}{1-\log_{2}((\alpha\beta)^{n}r_{0})}\cdot\frac{k_{n}}{k_{n+1}}$
$=\frac{-(a+2)}{-(a+b)}$
$=\frac{\log(4\alpha)}{\log(\alpha\beta)}>c$
as desired. Therefore, Alice wins. The losing condition is analogous, so we omit it.
For the proofs below, we will use the following
Lemma 5.4.
Let $S\subseteq\mathbb{N}$ be an infinite set of density zero. Let $d^{\pm}(x,0)=c$; then
$d^{\pm}\left(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0\right)\geq c.$
Proof.
For $x\in\mathbb{Z}[\frac{1}{2}]$ one has $d^{\pm}t(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0)=d^{\pm}(\sum\limits_{j\in S}\frac{1}{2^{j}},0)=1$, as addition of $x$ changes only finitely many digits. Now assume $x\notin\mathbb{Z}[\frac{1}{2}]$.
[…]
$d^{\pm}(x+\sum\limits_{j\in S_{2}}\frac{1}{2^{j}},0)={\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\sqcup S_{2}\}}{k}=$
${\lim\limits_{k\rightarrow\infty}}^{\pm}\Big{(}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\frac{\#\{1\leq i\leq k:\,i\in S_{2}\}}{k}\Big{)}=$
${\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-\lim\limits_{k\rightarrow\infty}\frac{\#\{1\leq i\leq k:\,i\in S_{2}\}}{k}=$
$={\lim\limits_{k\rightarrow\infty}}^{\pm}\frac{\#\{1\leq i\leq k:\,i\notin Q\}}{k}-0=d^{\pm}(x,0)=c.$
Now it remains to prove the statement only for $S$ such that $S\subseteq Q$. Let us define $R\subseteq\mathbb{N}$ by $\sum\limits_{j\in Q}\frac{1}{2^{j}}+\sum\limits_{j\in S}\frac{1}{2^{j}}=\sum\limits_{j\in R}\frac{1}{2^{j}}$. It is easy to check that:
  • Suppose $i\notin Q$ and $i+1\notin Q$. Then $i\notin R$;
  • Suppose $a\notin Q,\,a+b\notin Q$ and $i\in Q$ for all $a<i<a+b$. Then #$\#\{a\leq i<a+b:\,i\in R\}\leq b-1=\#\{a\leq i<a+b:\,i\in Q\}$.
This immediately implies that $d^{\pm}(x+\sum\limits_{j\in S}\frac{1}{2^{j}},0)\geq d^{\pm}(x,0)=c$ and completes the proof. ∎
Let us also recall simple properties of digit frequencies:
  • $d^{\pm}(2^{k}x,i)=d^{\pm}(2^{l}+x,i)$ for any integers $k,l$;
  • $d^{\pm}(x,i)+d^{\mp}(x,1-i)=1$;
  • If $x\notin\mathbb{Z}[\frac{1}{2}]$, then $d^{\pm}(x,i)=d^{\pm}(-x,1-i)$.
We now focus on the special case $c=\frac{1}{2}$. We will use the construction similar to the set $\mathscr{D}$ but connected to digit frequencies.
Let $\mathscr{F}=D^{-}_{\frac{1}{2}}\cup(\mathscr{D}(\xi)\cap\{x:\,d^{+}(x,0)\geq\frac{1}{2}\})$ where $\xi=\sum\limits_{i=1}^{\infty}\frac{1}{2^{2^{i}}}$. We want to show that the set $\mathscr{F}$ satisfies the conditions of Theorem 3.10 for $m=2$.
Proposition 5.5.
  1. (1)
    $2^{k}\mathscr{F}=2^{k}+\mathscr{F}=\mathscr{F}$ for any integer $k$;
  2. (2)
    $\mathscr{F}\sqcup(-\mathscr{F})\sqcup Q=\mathbb{R}$ where $Q\subseteq\mathbb{Q}$;
  3. (3)
    $\mathscr{F}+\xi\subseteq\mathscr{F}$.
Proof.
  1. (1)
    Follows from the similar properties of all the sets in construction of $\mathscr{F}$;
  2. (2)
    It’s easy to notice that
    =$-\mathscr{F}=\left\{x\notin\mathbb{Z}\left[\frac{1}{2}\right]:\,d^{+}(x,0)<\frac{1}{2}\right\}\cup\left((-\mathscr{D}(\xi))\cap\left\{x:\,d^{-}(x,0)\leq\frac{1}{2}\right\}\right),$
    which shows that $\mathscr{F}$ and $-\mathscr{F}$ are disjoint. If $x\notin\mathscr{F}\sqcup(-\mathscr{F})$, then either $x\in\mathbb{Z}[\frac{1}{2}]$ or $d^{-}(x,0)\leq\frac{1}{2}\leq d^{+}(x,0)$ and $x\notin\mathscr{D}\sqcup(-\mathscr{D})$, which implies $x\in\mathbb{Q}$.
  3. (3)
    By Lemma 5.4 all the sets in the construction of $\mathscr{F}$ satisfy this condition.
This immediately proves that
  • If $\alpha\geq\beta$, then the set $\mathscr{F}$ is $(\alpha,\beta)$-losing;
  • If $\beta>\alpha$ and $\log_{m}(\alpha\beta)\notin\mathbb{Q}$, then the set $\mathscr{F}$ is $(\alpha,\beta)$-winning.
Let’s also note that $\{x:\,d^{-}(x,0)>\frac{1}{2}\}\subseteq\mathscr{F}\subseteq\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}$, thus
Corollary 5.6.
  1. (1)
    If $\alpha\geq\beta$, then the set $D^{-}_{\frac{1}{2}}=\{x:\,d^{-}(x,0)>\frac{1}{2}\}$ is $(\alpha,\beta)$-losing;
  2. (2)
    If $\beta>\alpha$ and $\log_{2}(\alpha\beta)\notin\mathbb{Q}$, then the set $\{x:\,d^{+}(x,0)\geq\frac{1}{2}\}$ is $(\alpha,\beta)$-winning.
Conjecture 5.7.
We believe that $D(D^{-}_{c})=\{(\alpha,\beta):\,\log(\alpha)/\log(\alpha\beta)>c\}$.
6. B-Badly Approximable Numbers
The set of badly approximable numbers is an important object of study in Diophantine approximations. It is defined as follows:
$\operatorname{BA}:=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{n}\right|>\frac{c}{n^{2}}\text{ for some $c>0$ and all }m\in{\mathbb{Z}},\ n\in{\mathbb{N}}\right\}.$
It is known that the badly approximable numbers are dense and have measure zero. Interestingly enough, despite having measure zero, Schmidt showed that they are winning for the largest possible nontrivial bounds:
Theorem 6.1.
[S66] The set $\operatorname{BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$; that is, $D(\operatorname{BA})=\hat{D}$.
In a similar vein to the badly approximable numbers investigated by Schmidt, we define the set of B-Badly Approximable Numbers as
$\operatorname{B-BA}:=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for some $c>0$ and all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}$
and a closely related set
$\operatorname{B-BA}(c):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}.$
These definitions imply $\operatorname{B-BA}=\bigcup_{c>0}\operatorname{B-BA}(c)$. We note that the $\operatorname{B-BA}$ set has been previously studied in [N10] and [PS08].
The following statement follows from work of Schmidt:
Theorem 6.2.
[S66] The set $\operatorname{2-BA}$ is winning for all $(\alpha,\beta)\in\hat{D}$; that is, $D(\operatorname{2-BA})=\hat{D}$.
While both $\operatorname{2-BA}$ and $\operatorname{2-BA}(c)$ did not produce any new Schmidt diagrams, there is a simple change we can make that will produce both nontrivial winning and nontrivial losing zones. For any integer $B\geq 2$, define
$\operatorname{\widetilde{B-BA}}(c,N):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{B^{k}}\right|>\frac{c}{B^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\ \text{s.t.}\ k>N\right\}$
and
$\operatorname{\widetilde{B-BA}}(c):=\bigcup_{N\in{\mathbb{N}}}\operatorname{\widetilde{B-BA}}(c,N).$
Throughout this chapter, we define $\gamma_{A}=1-2\alpha+\alpha\beta$ and $\gamma_{B}=1-2\beta+\alpha\beta$. We begin results in this section with a useful general lemma:
Lemma 6.3.
Suppose Bob plays any interval $B_{t}$ and $\gamma_{A}>0$. Denote the center of $B_{t}$ by $c_{t}$. Then, Bob can always choose an interval $B_{t+1}$ such that $|c_{t+1}-c_{t}|\leq\frac{|B_{t}|}{2}\gamma_{A}$. As a corollary, Bob can guarantee that for any $m\in\mathbb{N}$, $|c_{t+m}-c_{t}|\leq\frac{|B_{t}|}{2(1-\alpha\beta)}\gamma_{A}$. Analogous results hold for Alice, with $\gamma_{A}$ replaced by $\gamma_{B}$.
Proof.
We prove Bob’s case; Alice’s case is analogous.
Let’s denote center of $B_{t}$ by $c_{t}$. It is clear that Bob can always choose such a $B_{t+1}$ that $|c_{r+1}-c_{r}|\leq\frac{|B_{t}|}{2}\gamma_{A}$, since no matter which side Alice picks her interval on, Bob will go the opposite way. Extending this logic, it follows that $|c_{m+r}-(a_{0}+k\delta)|\leq\frac{\rho(\alpha\beta)^{m}}{2}\gamma_{A}(1+(\alpha\beta)+(\alpha\beta)^{2}+\ldots)=\frac{\rho(\alpha\beta)^{m}}{2(1-\alpha\beta)}\gamma_{A}$, which completes the proof.
The following theorems provide the winning and losing conditions for $\operatorname{\widetilde{B-BA}}$ sets.
Theorem 6.4.
If
$\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma_{A}}>1,$
then for any $B$, the set $\operatorname{\widetilde{B-BA}}(c)$ is $(\alpha,\beta)$-losing.
Proof.
One has
$\operatorname{\widetilde{B-BA}}(c)^{c}=\left\{x:\,\,\forall N\,\,\exists\,n>N\,\,\text{and}\,\,k\,\in\,\mathbb{Z}:\,\,\left|x-\frac{k}{B^{n}}\right|<\frac{c}{B^{n}}\right\}=$
$=\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left[\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right]\right)\subseteq\bigcap\limits_{N=1}^{\infty}\bigcup\limits_{n=N}^{\infty}\left(\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right)\right),$
and let $J_{n}=\bigcup\limits_{k\in\mathbb{Z}}\left(\frac{k}{B^{n}}-\frac{c}{B^{n}},\frac{k}{B^{n}}+\frac{c}{B^{n}}\right)$. Fix $\varepsilon>0$ such that
$\varepsilon=\frac{2c(1-\beta)(1-\alpha\beta)}{\beta\gamma_{A}}-1.$
Bob will play in the following way:
  • On his first turn, Bob will play an interval of length
    $\frac{1}{B}\left(\frac{2c(1-\alpha\beta)}{\gamma_{A}}\right)$
    centered at $1/B$. By Lemma 1.2, Bob can play in such a way that in finitely many moves, $B_{s}\subset J_{1}$.
  • On all subsequent turns, Bob plays arbitrarily until some turn $t$ such that
    $\log_{B}\left(\frac{1}{1-\beta}\right)$$\leq\left\{\left(t\cdot\log_{B}\alpha\beta\right)+\log_{B}|A_{0}|\right\}$ (6.1)
    $\leq\log_{B}\left(\frac{1}{1-\beta}\right)+\log_{B}(1+\varepsilon),$
    there exists $N\in\mathbb{N}$ such that
    $\left(\frac{1}{1-\beta}\right)\leq B^{N}|A_{t}|\leq\left(\frac{1}{1-\beta}\right)(1+\varepsilon)=\left(\frac{2c(1-\alpha\beta)}{\beta\gamma_{A}}\right),$ (6.2)
    and since $|A_{t}|(1-\beta)>1/B^{N}$, Bob can play his next interval centered at $m/B^{N}$ for some $m\in\mathbb{Z}$. Furthermore, by Lemma 1.2, Bob can play in such a way that in finitely many moves, $B_{t+s}\subset J_{n}$. By the Equidistribution Theorem, there exists an infinite number of possible $t$ that satisfy inequality 6.2. Therefore, Bob can do this an infinite number of times, and Alice loses.
In the special case of the $\operatorname{\widetilde{2-BA}}$ set, we have the following theorem.
Theorem 6.5.
The set $\operatorname{\widetilde{2-BA}}(c)$ is winning for $\alpha<\frac{1}{3}$, $\beta>\frac{9c}{1-3\alpha}$, and $c<1/6$.
Proof.
The winning strategy for Alice is as follows: suppose Alice is given an interval $B_{i}$ of length $l_{i}$. Pick $k_{i}$ such that
$\frac{1}{2^{k_{i}}}<l_{i}\leq\frac{1}{2^{k_{i}-1}}$
or equivalently,
$\frac{l_{i}}{2}\leq\frac{1}{2^{k_{i}}}<l_{i}.$
Let $W_{i}$ denote the set of all intervals of length $\alpha l_{i}$ and center $c_{i}$ of the form $(m\pm\frac{1}{3})2^{-k_{i}}$, where $m\in{\mathbb{Z}}$. Then, Alice will select an interval from $W_{i}$. This is always possible because the maximum distance between the two endpoints of sets is
$\left|c_{i}-c^{\prime}_{i}\right|+\alpha l_{i}=\frac{1}{3\cdot 2^{k_{i}-1}}+\alpha l_{i}<\frac{2l_{i}}{3}+\frac{l_{i}}{3}=l_{i}$
and thus any interval of length $l_{i}$ must contain at least one of the intervals in $W_{i}$.
Next, define $q_{i}$ to be the smallest integer such that $m/2^{q_{i}}\in A_{i}$ for some $m\in{\mathbb{Z}}$. Note that
$\left|\frac{m}{2^{q_{i}}}-c_{i}\right|=\frac{1}{3\cdot 2^{q_{i}}}<\frac{\alpha l_{i}}{2}<\frac{1}{3\cdot 2^{q_{i}-1}}.$
We wish to show that for all $i\geq 1$ and $x\in A_{i}$,
$\left|x-\frac{m}{2^{n}}\right|>\frac{c}{2^{n}}\ \text{for all}\ n\leq q_{i}-2.$
by induction on $i$. Suppose that the statement holds for $i$. Then, Bob picks an interval $B_{i+1}$ with length $l_{i+1}=\alpha\beta l_{i}$. Now, note that for all $x\in A_{i+1}$ and $q_{i}\leq n\leq k_{i+1}$,
$\left|x-\frac{m}{2^{n}}\right|\geq\left|x-\frac{m^{\prime}}{2^{k_{i+1}}}\right|$ $>\frac{1}{3\cdot 2^{k_{i+1}}}-\frac{\alpha l_{i+1}}{2}$
$\geq\frac{l_{i+1}}{6}-\frac{\alpha l_{i+1}}{2}$
$=\alpha\beta l_{i}\left(\frac{1-3\alpha}{6}\right)$
$>\alpha l_{i}\left(\frac{9c}{1-3\alpha}\right)\left(\frac{1-3\alpha}{6}\right)$
$=\frac{3}{2}\alpha l_{i}c$
$>\frac{c}{2^{q_{i}}}\geq\frac{c}{2^{n}}.$
To take care of $k_{i+1}+1\leq n\leq q_{i+1}-2$, simply note that for all $x\in A_{i+1}$ and $c<1/6$,
$\left|x-\frac{m}{2^{n}}\right|$ $>\frac{1}{3\cdot 2^{n}}-\frac{\alpha l_{i+1}}{2}>\frac{1}{3\cdot 2^{n}}-\frac{1}{3\cdot 2^{q_{i+1}-1}}>\frac{c}{2^{n}}.$
Finally, for $n=q_{i}-1$, we split into cases based on $A_{i}$, the prior move.
  • Case I.
    If
    $\frac{\alpha l_{i}}{2}<\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{c}{2^{q_{i}-1}},$
    then for all $x\in A_{i}$,
    $\left|x-\frac{m}{2^{q_{i}-1}}\right|$ $\geq\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{\alpha l_{i}}{2}$
    $>\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{1}{3\cdot 2^{q_{i}-1}}+\frac{c}{2^{q_{i}-1}}$
    $=\frac{c}{2^{q_{i}-1}}.$
  • Case II.
    Assume
    $\frac{\alpha l_{i}}{2}\geq\frac{1}{3\cdot 2^{q_{i}-1}}-\frac{c}{2^{q_{i}-1}}.$
    Then, if
    $B_{i+1}\cap\left[\frac{m-c}{2^{q_{i}-1}},\frac{m+c}{2^{q_{i}-1}}\right]=\emptyset,$
    it is clear that Alice does not have to worry about $n=q_{i}-1$, since for all $x\in A_{i+1}\subset B_{i+1}$, $|x-m/2^{q_{i}-1}|>c/2^{q_{i}+1}$. Thus, for the remainder of this case, assume $B_{i+1}$ intersects the interval. Without loss of generality, assume that $m/2^{q_{i}-1}>x$ for all $x\in B_{i+1}$ (that is, the fraction lies to the right of the interval $B_{i+1}$). This is possible because the mapping $x\mapsto-x$ gives us the other case. Alice will then choose an interval $A_{i+1}\in W_{i+1}$ centered at $c_{i+1}$ such that
    $c_{i+1}\leq\frac{m}{2^{q_{i}-1}}-\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right),$
    that is, Alice cannot choose the closest interval to $\frac{m}{2^{q_{i}-1}}$. This is always possible because the second closest interval is within $l_{i+1}$ of $\frac{m}{2^{q_{i}-1}}$; that is, for $x\in A_{i+1}$,
    $\left|x-\frac{m}{2^{q_{i}-1}}\right|<\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right)+\frac{\alpha l_{i+1}}{2}<\frac{2}{3}l_{i+1}+\frac{1}{6}l_{i+1}<l_{i+1}$
    and thus whenever Alice is able to pick the closest interval, Alice is also able to pick the second closest interval. Furthermore, for $x\in A_{i+1}$,
    $\left|x-\frac{m}{2^{q_{i}-1}}\right|$ $>\frac{2}{3}\left(\frac{1}{2^{k_{i+1}}}\right)-\frac{\alpha l_{i+1}}{2}$
    $>\frac{l_{i+1}}{3}-\alpha l_{i+1}$
    $=\alpha\beta l_{i}\left(\frac{1-3\alpha}{3}\right)$
    $>3\alpha l_{i}c$
    $>\frac{c}{2^{q_{i}-1}}$
    proving this case and completing the induction.
Thus, $x_{\infty}\in\operatorname{\widetilde{2-BA}}(c)$ and Alice wins.
A special case of winning and losing zones given by Theorem 6.4 and Theorem 6.5 can be seen on Figure 5.
We also consider the special case of $\widetilde{\text{3-BA}}(1/6)$, defined as
$\widetilde{\text{3-BA}}(1/6):=\left\{x\in{\mathbb{R}}:\left|x-\frac{m}{3^{k}}\right|\geq\frac{1}{6\cdot 3^{k}}\text{ for all }m\in{\mathbb{Z}},\ k\in{\mathbb{N}}\right\}.$
There, we have some additional bounds:
Theorem 6.6.
If
$\beta\gamma_{A}<\gamma_{B}$
and
$\frac{1-\alpha\beta}{\alpha\gamma_{B}}\geq\frac{9\beta}{1-\beta}$
and
$\alpha\beta>\frac{1}{7},$
then the set $\widetilde{\text{3-BA}}(1/6)$ is $(\alpha,\beta)$-losing.
Proof.
Let
$2\varepsilon={\gamma_{B}}-\beta\gamma_{A}.$
Then, Bob will play in the following way: on his first turn, he plays an interval of length
$|B_{0}|=\frac{1-\alpha\beta}{\alpha\gamma_{B}}.$
Then, he will play arbitrarily until some turn $t$, where
$\frac{1-\alpha\beta}{\alpha\gamma_{B}}\cdot\frac{1}{3^{n+2}}\leq|B_{t}|<(1+\varepsilon)\cdot\frac{1-\alpha\beta}{\alpha\gamma_{B}}\cdot\frac{1}{3^{n+2}}<\frac{1-\alpha\beta}{\alpha\beta\gamma_{A}}\cdot\frac{1}{3^{n+2}}$
for some $n\in\mathbb{N}$. Then, on turn $t$, he will center his ball at a $c_{t}$ such that
$c_{t}=\frac{m}{3^{n}}$
for some $m\in\mathbb{Z}$. On his next move, Bob is always able to play in one of the following two ways:
  • Bob will center his next interval such that
    $\left|c_{t+1}-\frac{m}{3^{n}}\right|<\frac{1}{6\cdot 3^{n}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ (6.3)
    and Bob will try to minimize $|c_{t+1}-\frac{m}{3^{n}}|$, or
  • Bob will center his next interval such that
    $\left|c_{t+1}-\left(\frac{m}{3^{n}}\pm\frac{1}{3^{n+1}}\right)\right|<\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ (6.4)
    and Bob will try to minimize $|c_{t+1}-\frac{m}{3^{n}}\mp\frac{1}{3^{n}+1}|$.
We show that this is always possible. First, we show that Bob can play his strategy on turn $t$. The length condition is equivalent to
$0\leq\left\{\log_{3}\frac{|B_{0}|\alpha\gamma_{B}}{1-\alpha\beta}+t\log_{3}\alpha\beta\right\}<\log_{3}(1+\varepsilon)$
and by the Equidistribution Theorem, there are infinitely many values of $t$ that satisfy the inequality. Then, we have
$|B_{t}|\geq\frac{1-\alpha\beta}{\alpha\beta\gamma_{A}}\cdot\frac{1}{3^{n+2}}\geq\frac{\beta}{1-\beta}\cdot\frac{1}{3^{n}}$
so $(1-\beta)|A_{t-1}|\geq 1/3^{n}$ and therefore Bob can always play an interval of the form $m/3^{n}$ on turn $t$.
Now, we show Bob can play his strategy on turn $t+1$. First, we note that since $\alpha\beta>1/7$, we have
$\frac{1}{6\alpha\beta}$ $<\frac{7}{6}$
$\frac{|B_{t}|\gamma_{A}}{2(1-\alpha\beta)}<\frac{1}{2\alpha\beta}\cdot\frac{1}{3^{n+2}}$ $<\frac{7}{6\cdot 3^{n+1}}$
$\frac{|B_{t}|\gamma_{A}}{2}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$ $<\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}$
$\frac{m}{3^{n}}+\frac{|B_{t}|\gamma_{A}}{2}$ $<\frac{m}{3^{n}}+\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)},$
which implies that Bob is able to ensure that $c_{t+1}<\frac{m}{3^{n}}+\frac{1}{3^{n+1}}+\frac{1}{6\cdot 3^{n+1}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$. By similar logic, Bob can ensure that $c_{t+1}>\frac{m}{3^{n}}-\frac{1}{3^{n+1}}-\frac{1}{6\cdot 3^{n+1}}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}$. Furthermore, we note that
$\frac{|B_{t}|\alpha\gamma_{B}}{1-\alpha\beta}$ $>\frac{1}{3^{n+2}}$
$|B_{t}|\alpha\left(1-\beta-\frac{\beta\gamma_{A}}{1-\alpha\beta}\right)$ $>\frac{1}{3^{n+2}}$
$\frac{1}{3^{n+2}}+\frac{|B_{t}|\alpha\beta\gamma_{A}}{1-\alpha\beta}$ $<(1-\beta)\alpha|B_{t}|$
$\left|\left(\frac{m}{3^{n}}\pm\left(\frac{1-\frac{1}{6}}{3^{n+1}}+\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}\right)\right)-\left(\frac{m}{3^{n}}\pm\left(\frac{1}{6\cdot 3^{n}}-\frac{|B_{t+1}|\gamma_{A}}{2(1-\alpha\beta)}\right)\right)\right|$ $<(1-\beta)\alpha|B_{t}|,$
which implies that Bob is always able to satisfy either 6.3 or 6.4.
Finally, we show this strategy is winning. This follows from Lemma 6.3: if 6.3 is true, then in finitely many $k$ turns, $|B_{t+k}|\subset(\frac{m-1/6}{3^{n}},\frac{m+1/6}{3^{n}})$. Since Bob can play this way for infinitely many sets of two consecutive turns, Bob wins. ∎
Conjecture 6.7.
If
$\beta\gamma_{A}\geq\gamma_{B},$
then the set $\widetilde{\text{3-BA}}(1/6)$ is $(\alpha,\beta)$-winning. If not, then it is $(\alpha,\beta)$-losing.
Acknowledgements
We are very grateful to Prof. Dmitry Kleinbock for the project proposal and his valuable suggestions in research paths. We would also like to thank Prof. Pavel Etingof, Dr. Slava Gerovitch, Dr. Felix Gotti, Dr. Tanya Khovanova, and all the PRIMES organizers for providing the opportunity to conduct math research through this program.